OpenStudy (anonymous):

Dante spent $11.25 on 3 new and 4 old comic books, and Samantha spent $15.75 0n 10 old and 3 new ones. If comics of one type are sold at the same price, what is the price in dollars of a new comic book?

OpenStudy (anonymous):

Make new comics = x Old comics = y 3x + 4y = 1125 (1) 3x + 10y = 1575 (2) Subtract (1) from (2) 6y = 450 y = 450/6 y = 75

OpenStudy (anonymous):

@jojor_124